# Designing Multi-Robot Ground Video Sensemaking with Public Safety Professionals

### v1

@@ Abstract.
Videos from fleets of ground robots can advance public safety by providing scalable situational awareness and reducing professionals’ burden. Yet little is known about how to design and integrate multi-robot videos into public safety workflows. Collaborating with six police agencies, we examined how such videos could be made practical. In Study 1, we presented the first testbed for multi-robot ground video sensemaking. The testbed includes 38 events-of-interest (EoI) relevant to public safety, a dataset of 20 robot patrol videos (10 day/night pairs) covering EoI types, and 6 design requirements aimed at improving current video sensemaking practices. In Study 2, we built MRVS, a tool that augments multi-robot patrol video streams with a prompt-engineered video understanding model. Participants reported reduced manual workload and greater confidence with LLM-based explanations, while noting concerns about false alarms and privacy. We conclude with implications for designing future multi-robot video sensemaking tools. The testbed is available at https://github.com/Puqi7/MRVS_VideoSensemaking.
@@ 1. Introduction
While ground robotics could play a transformative role in high-stakes public-safety operations, we still lack a clear understanding of how these systems should be designed, built, and evaluated in ways that reflect public-safety professionals’ real practices, constraints, and decision-making needs. To mitigate this gap, this work first presents a testbed environment for multi-robot video sensemaking , co-designed with experienced public-safety professionals, that future researchers can build upon (Study 1, see Fig. 1, S1). This testbed addresses the current gap by providing a first-of-its-kind design and evaluation environment capable of supporting multi-robot, multi-video workflows in realistic public-safety operational settings. Next, this work assesses the testbed’s practicality with public safety professionals through the development of an interactive  Multi-Robot Video Sensemaking System, (MRVS; Study 2, see Fig. 1, S2), conducted in collaboration with six local and state public safety agencies. MRVS is the first interactive system that operationalizes multi-robot video sensemaking for public safety by tightly aligning a multimodal AI back-end and a human-centered front-end with public safety professionals’ procedures, concerns, and adoption constraints.
@@ 3.2.1. Methodology
Step 1. EoIs Preparation: We first constructed a candidate set of EoIs balancing operational realism, practical relevance, and reduced participant burden through a three-step process: (1) Data collection: We assembled an initial pool of candidate records from two sources: (a) 13,234 daily crime logs collected over three years across three universities in one of the U.S. states and (b) ten public research video datasets on anomaly detection and crime video analysis (Ma et al., 2015; Ramachandra and Jones, 2020; Qian et al., 2025; Wu et al., 2020; Wang et al., 2020a; Zhang et al., 2016; Acsintoae et al., 2022; Pranav et al., 2020; abnormal2013lu; Rodrigues et al., 2020). For the daily crime logs, we extracted the existing fields Nature of Case, Brief Description, and Offenses. For the video datasets, we collected available event types and descriptions; when unavailable, two authors independently reviewed videos, drafted descriptions, and resolved differences through consensus coding (Strauss, 1987) to complete our initial record archive. (2) Filtering records: Using descriptions from both sources, we removed records that (a) could not be visually detected from robot-captured footage or (b) would be impractical to simulate (e.g., sexual offenses). (3) Classification: With input from our public-safety co-author, we reviewed all remaining records and defined event types aligned with U.S. public-safety standards. Category naming drew from the Clery Act (Center, ), a legal standard for campus-crime classification, and event types used in existing abnormal-event video datasets. To categorize the 40 candidate events of interest (EoIs), five authors conducted an in-person collaborative affinity diagramming session (Lucero, 2015). Authors iteratively clustered EoIs and developed high-level themes through synchronous discussion, informed by a coauthor’s public safety domain expertise. Each EoI was assigned only after unanimous consensus, followed by secondary review to ensure themes were mutually exclusive and differentiated. This produced six categories: (a) Property and Environmental Incidents, (b) Public Order Disturbances, (c) Vehicle and Mobility Incidents, (d) Suspicious or Unusual Behavior, (e) High-Risk Threats, and (f) Miscellaneous Activities. Finally, taxonomy was refined with concise definitions and representative visual cues to ground the survey. Each category was accompanied by concise definitions and representative visual cues derived during data collection, with example images sourced from public materials, establishing a structured foundation for the survey.
@@ 19 Road Accidents Urgent 6 Civil 38 Propping Doors Open Advisory 1 Civil
Table 2. Event Classification with Occurrence Count in testbed environment
@@ 4. Video Dataset
We present a first-of-its-kind video dataset [^1] Link to Dataset: https://puqi7.github.io/MRVS_VideoSensemaking/. captured by a patrolling ground robot to establish a testbed environment for multi-robot video sensemaking. Unlike existing anomaly video datasets, ours is grounded in co-defined EoIs with public safety professionals, enacted in real-world environments, and recorded under police-guided patrol scripts specifying locations and event types. This dataset offers a realistic benchmark for evaluating future systems leveraging robot-captured video in public safety. Representative scripted EoIs are shown in Figure 2.
Grounded in the refined EoI types in Table 3.3.2, we captured 156 scripted events across 10 campus zones and 10 day/night sessions with 22 student actors, resulting in 20 non-stop patrol videos around 30 minutes each, containing 833 people and 1,537 vehicles. For video capture, we utilized a Frodo Zero ground robot (Frodobots, ) equipped with a GoPro Hero 11 (GoPro, ), recording 4K-resolution videos without audio during video taking. Each video contains GPS coordinates and timestamps. Our video production process is as follows.
Step 1. Preparation Before the main video production, we prepared four elements to prioritize ecological validity, ensuring footage appeared natural, realistic, and reusable for future studies. Specifically, we focused on defining plausible patrol routes, creating naturalistic acting guidelines, conducting dry-run pilots, and securing public safety professional oversight. First, robot patrol routes were designed to cover campus “hot spots” identified by the eighth author with public-safety domain expertise, remain fully wheel-accessible, and avoid spatial overlap. Second, the acting scripts specified naturalistic behavior, such as maintaining a normal walking pace, avoiding exaggerated glances toward the robot, and performing each EoI at varying distances, locations, and occlusion levels to reflect real-world visual complexity. We balanced the distribution of EoI types and ensured that each EoI was represented at least once in the dataset. Third, we conducted dry-run recordings using two scripted routes and seven actors, reviewing two 30-minute pilot videos to tune robot speed, route length, and ensure actor behaviors appeared natural. Finally, the eighth author confirmed these behaviors aligned with actual campus incidents, leading us to refine both patrol routes and acting scripts before final production. We note that some abnormal events cannot be naturally captured, so we followed prior work (Lu et al., 2025; de Souza et al., 2017) using actor-performed scenarios to ensure systematic coverage of rare and critical events.
@@
Step 3. Post Production: Anonymization and Distribution All identifiable faces and license plates were blurred using ORB-HD Deface (ORB-HD, 2025) for facial anonymization and EgoBlur (Raina et al., 2023), which is based on Faster R-CNN (Ren et al., 2016), for license-plate anonymization. This produces a research-safe foundation for evaluating timeline-based filtering, real-time alerting, descriptor-based search, and event detection. The dataset contains video only, with no audio, to avoid potential First Amendment concerns (Freedom Forum, ). A research-safe version will be shared publicly, including (a) anonymized video files (with all faces and license plates blurred) and (b) metadata describing EoI types, timestamps, and GPS information.
@@ 5.2.1. MLLM-based Video Sensemaking (DR1-DR3)
Specifically, our approach builds upon the capabilities of Gemini 2.0 Flash (Team et al., 2024) as a baseline based on its leading performance on established video understanding benchmarks (Wu et al., 2024a; Wang et al., 2025; Salehi et al., 2024) and enhances it with multi-modal reasoning and spatial-temporal localization. The model was configured with a temperature of 1.0 to balance response creativity with output consistency and reliability. We guide attention to objects of interest with point prompting (Dai et al., 2024; Dai and Yang, 2024). Specifically, we preprocess the video streams using the BoxMOT (Broström, ) framework, which integrates real-time multi-object tracking with YOLO (Wang et al., 2024a) based detectors. Each frame is enriched with bounding boxes $\mathcal{B}=\{b_{i}=(x_{1,i},y_{1,i},x_{2,i},y_{2,i})\}_{i=1}^{N}$ and class labels $\mathcal{C}=\{c_{i}\}_{i=1}^{N}$ for detected entities. These annotated frames form an object-aware video stream $\mathcal{V}^{*}$, which is then processed by a multimodal large language model (MLLM) using both role prompting and temporal segmentation. Inspired by (Kong et al., 2024), we introduce a system-level persona prompt, defining the model as a specialized surveillance agent (VSAI-9000) with domain-specific analysis protocols. This role-based formulation enhances zero-shot reasoning and structured event classification. The input video is temporally segmented into overlapping intervals $\mathcal{S}=\{s_{j}\}_{j=1}^{n}$ (e.g., 10 seconds each), where consecutive segments share partial temporal overlap to ensure continuous coverage of key activities and smooth transition detection. The 10 second interval duration was determined through pilot testing with 10 videos, balancing sufficient temporal context for meaningful event identification against computational efficiency and the model’s context window limitations. For each interval, the pipeline (Fig. 7) returns the following: the event type $e\in\mathcal{E}$, a natural language description $\mathcal{D}(e)$, an explanatory rationale $\mathcal{R}(e)$, a confidence level $l(e)\in\{\textit{high},\textit{medium},\textit{low}\}$, and representative frame(s) $\mathcal{K}(e)$ selected via keyframe extraction. A keyframe is selected as the frame with the highest count of distinct objects. This frame serves as the clearest snapshot of the interval. The model is instructed to state its confidence as one of three discrete levels, high, medium, or low, based on how certain it is about the inferred event type and description. This confidence level is passed unchanged to the front-end (e.g. shown as “High” in Fig. 3) and used as an ordinal indicator of reliability. This design follows recent work showing that large language models can express reasonably calibrated uncertainty in natural language without exposing logits or explicit probability scores (Lin et al., 2022; Kadavath et al., 2022).
@@ 5.2.2. Descriptor-based Search (DR5)
Objects are extracted from videos using BoxMOT (Broström, ), combining YOLO-based object detectors (Redmon et al., 2016) with DeepOCSort (Maggiolino et al., 2023) for multi-object tracking. LLAVA 1.6 (Liu et al., 2024a) verifies detection semantic correctness through binary questions, discarding negatives. For cars and persons, descriptor-based attributes are extracted through structured queries, with cropped first-appearance images stored for validation. Predicted attributes are re-evaluated on saved crops, and rejected outputs are placed in the “other” category.
@@ 7.1.1. Testbed environment
  • The EoIs Types (see Table 2) were informed by prior anomaly detection video datasets that define “anomaly” from a computer vision perspective as statistical rarity (Mahadevan et al., 2010), researcher-defined categories (Ramachandra and Jones, 2020), or labeled from online surveillance footage (Wu et al., 2020), valuable for benchmarking but not directly actionable in public safety practice. We drew EoIs from real campus crime logs (following Clery Act (Center, )) on the foundation of existing anomaly datasets (Ma et al., 2015; Ramachandra and Jones, 2020; Qian et al., 2025; Wu et al., 2020; Wang et al., 2020a; Zhang et al., 2016; Acsintoae et al., 2022; Pranav et al., 2020; Lu et al., 2013; Rodrigues et al., 2020), then refined with frontline professionals to yield operationally meaningful, context-rich categories (e.g., “Incident Exposure”). This approach grounds “what to detect” in professionals’ operational realities, translating their priorities into actionable EoIs for future system and model design. However, our EoIs remain campus-centric, reflecting a limited jurisdictional perspective and excluding non-visual or hard-to-simulate incidents. We therefore position it not as a universal ontology, but as a reusable approach: MRVS-like systems should expose EoIs as configurable, locally governed “recipes” rather than fixed, one-size-fits-all classes. The EoI types serve as a reusable template that other jurisdictions can adapt, extend to additional modalities (e.g., audio, text), and use to drive MRVS-like systems and upstream workflows (e.g., dispatch (Liu et al., 2023; Terrell et al., 2004; Munro et al., 2025)).
@@ Acknowledgements.
This work was supported by the National Science Foundation (NSF 2128867, 2350352), the Army Research Office (W911NF2320004, W911NF2520011), Google DeepMind (GDM), Clearpath Robotics, FrodoBots Lab, Raytheon Technologies (RTX), Tangenta, the Mason Innovation Exchange (MIX), and Walmart. We thank the George Mason University undergraduate students for their help with video dataset collection. We are grateful to the Northern Virginia police departments for their support and domain guidance, especially Detective Cooper Knight, Major Bull Hudson, Lieutenant Mohammed Tabibi, and Captain Andrew E. Hawkins, as well as other colleagues who provided feedback and guidance. We used generative AI tools for proofreading and editing under full human supervision, in line with the GAIDeT taxonomy (2025) (Suchikova et al., 2025). The authors are solely responsible for the final manuscript and outcomes.

### v2

@@ Abstract.
Videos from fleets of ground robots can advance public safety by providing scalable situational awareness and reducing professionals’ burden. Yet little is known about how to design and integrate multi-robot videos into public safety workflows. Collaborating with six police agencies, we examined how such videos could be made practical. In Study 1, we present the first testbed for multi-robot ground video sensemaking. The testbed includes 38 events of interest relevant to public safety, a dataset of 20 robot patrol videos (10 day/night pairs) covering EoI types, and 6 design requirements aimed at improving current video sensemaking practices. In Study 2, we built MRVS, a tool that augments multi-robot patrol video streams with a prompt-engineered video understanding model. Participants reported reduced manual workload and greater confidence with LLM-based explanations, while noting concerns about false alarms and privacy. We conclude with implications for designing future multi-robot video sensemaking tools.
@@ 1. Introduction
While ground robotics could play a transformative role in high-stakes public-safety operations, we still lack a clear understanding of how these systems should be designed, built, and evaluated in ways that reflect public-safety professionals’ real practices, constraints, and decision-making needs. To mitigate this gap, this work first presents a testbed environment [^1] Link to Testbed: https://github.com/Puqi7/MRVS_VideoSensemaking for multi-robot video sensemaking , co-designed with experienced public-safety professionals, that future researchers can build upon (Study 1, see Fig. 1, S1). This testbed addresses the current gap by providing a first-of-its-kind design and evaluation environment capable of supporting multi-robot, multi-video workflows in realistic public-safety operational settings. Next, this work assesses the testbed’s practicality with public safety professionals through the development of an interactive  Multi-Robot Video Sensemaking System, (MRVS; Study 2, see Fig. 1, S2), conducted in collaboration with six local and state public safety agencies. MRVS is the first interactive system that operationalizes multi-robot video sensemaking for public safety by tightly aligning a multimodal AI back-end and a human-centered front-end with public safety professionals’ procedures, concerns, and adoption constraints.
@@ 3.2.1. Methodology
Step 1. EoIs Preparation: We first constructed a candidate set of EoIs balancing operational realism, practical relevance, and reduced participant burden through a three-step process: (1) Data collection: We assembled an initial pool of candidate records from two sources: (a) 13,234 daily crime logs collected over three years across three universities in one of the U.S. states and (b) ten public research video datasets on anomaly detection and crime video analysis (Ma et al., 2015; Ramachandra and Jones, 2020; Qian et al., 2025; Wu et al., 2020; Wang et al., 2020a; Zhang et al., 2016; Acsintoae et al., 2022; Pranav et al., 2020; abnormal2013lu; Rodrigues et al., 2020). For the daily crime logs, we extracted the existing fields Nature of Case, Brief Description, and Offenses. For the video datasets, we collected available event types and descriptions; when unavailable, two authors independently reviewed videos, drafted descriptions, and resolved differences through consensus coding (Strauss, 1987) to complete our initial record archive. (2) Filtering records: Using descriptions from both sources, we removed records that (a) could not be visually detected from robot-captured footage or (b) would be impractical to simulate (e.g., sexual offenses). (3) Classification: With input from our public-safety co-author, we reviewed all remaining records and defined event types aligned with U.S. public-safety standards. Category naming drew from the Clery Act (Center, 2026), a legal standard for campus-crime classification, and event types used in existing abnormal-event video datasets. To categorize the 40 candidate events of interest, five authors conducted an in-person collaborative affinity diagramming session (Lucero, 2015). Authors iteratively clustered EoIs and developed high-level themes through synchronous discussion, informed by a coauthor’s public safety domain expertise. Each EoI was assigned only after unanimous consensus, followed by secondary review to ensure themes were mutually exclusive and differentiated. This produced six categories: (a) Property and Environmental Incidents, (b) Public Order Disturbances, (c) Vehicle and Mobility Incidents, (d) Suspicious or Unusual Behavior, (e) High-Risk Threats, and (f) Miscellaneous Activities. Finally, taxonomy was refined with concise definitions and representative visual cues to ground the survey. Each category was accompanied by concise definitions and representative visual cues derived during data collection, with example images sourced from public materials, establishing a structured foundation for the survey.
@@ 19 Road Accidents Urgent 6 Civil 38 Propping Doors Open Advisory 1 Civil
Table 2. Events Classification with Occurrence Count in the testbed environment
@@ 4. Video Dataset
We present a first-of-its-kind video dataset [^2] Link to Dataset: https://huggingface.co/datasets/Puqi7/MRVS_anomaly_long_video_dataset captured by a patrolling ground robot to establish a testbed environment for multi-robot video sensemaking. Unlike existing anomaly video datasets, ours is grounded in co-defined EoIs with public safety professionals, enacted in real-world environments, and recorded under police-guided patrol scripts specifying locations and event types. This dataset offers a realistic benchmark for evaluating future systems leveraging robot-captured video in public safety. Representative scripted EoIs are shown in Fig. 2.
Grounded in the refined EoI types in Table 3.3.2, we captured 156 scripted events across 10 campus zones and 10 day/night sessions with 22 student actors, resulting in 20 non-stop patrol videos around 30 minutes each, containing 833 people and 1,537 vehicles. For video capture, we utilized a Frodo Zero ground robot (Frodobots, 2026) equipped with a GoPro Hero 11 (GoPro, 2026), recording 4K-resolution videos without audio during video taking. Each video contains GPS coordinates and timestamps. Our video production process is as follows.
Step 1. Preparation Before the main video production, we prepared four elements to prioritize ecological validity, ensuring footage appeared natural, realistic, and reusable for future studies. Specifically, we focused on defining plausible patrol routes, creating naturalistic acting guidelines, conducting dry-run pilots, and securing public safety professional oversight. First, robot patrol routes were designed to cover campus “hot spots” identified by the eighth author with public-safety domain expertise, remain fully wheel-accessible, and avoid spatial overlap. Second, the acting scripts specified naturalistic behavior, such as maintaining a normal walking pace, avoiding exaggerated glances toward the robot, and performing each EoI at varying distances, locations, and occlusion levels to reflect real-world visual complexity. We balanced the distribution of EoI types and ensured that each EoI was represented at least once in the dataset. Third, we conducted dry-run recordings using two scripted routes and seven actors, reviewing two 30-minute pilot videos to tune robot speed, route length, and ensure actor behaviors appeared natural. Finally, the eighth author confirmed these behaviors aligned with actual campus incidents, leading us to refine both patrol routes and acting scripts before final production.
@@
We note that some abnormal events cannot be naturally captured, so we followed prior work (Lu et al., 2025; de Souza et al., 2017) using actor-performed scenarios to ensure systematic coverage of rare and critical events.
@@
Step 3. Post Production: Anonymization and Distribution All identifiable faces and license plates were blurred using ORB-HD Deface (ORB-HD, 2025) for facial anonymization and EgoBlur (Raina et al., 2023), which is based on Faster R-CNN (Ren et al., 2016), for license-plate anonymization. This produces a research-safe foundation for evaluating timeline-based filtering, real-time alerting, descriptor-based search, and event detection. The dataset contains video only, with no audio, to avoid potential First Amendment concerns (Freedom Forum, 2026). A research-safe version will be shared publicly, including (a) anonymized video files (with all faces and license plates blurred) and (b) metadata describing EoI types, timestamps, and GPS information.
@@ 5.2.1. MLLM-based Video Sensemaking (DR1-DR3)
Specifically, our approach builds upon the capabilities of Gemini 2.0 Flash (Team et al., 2024) as a baseline based on its leading performance on established video understanding benchmarks (Wu et al., 2024a; Wang et al., 2025; Salehi et al., 2024) and enhances it with multi-modal reasoning and spatial-temporal localization. The model was configured with a temperature of 1.0 to balance response creativity with output consistency and reliability. We guide attention to objects of interest with point prompting (Dai et al., 2024; Dai and Yang, 2024). Specifically, we preprocess the video streams using the BoxMOT (Broström, 2023) framework, which integrates real-time multi-object tracking with YOLO (Wang et al., 2024a) based detectors. Each frame is enriched with bounding boxes $\mathcal{B}=\{b_{i}=(x_{1,i},y_{1,i},x_{2,i},y_{2,i})\}_{i=1}^{N}$ and class labels $\mathcal{C}=\{c_{i}\}_{i=1}^{N}$ for detected entities. These annotated frames form an object-aware video stream $\mathcal{V}^{*}$, which is then processed by a multimodal large language model (MLLM) using both role prompting and temporal segmentation. Inspired by (Kong et al., 2024), we introduce a system-level persona prompt, defining the model as a specialized surveillance agent (VSAI-9000) with domain-specific analysis protocols. This role-based formulation enhances zero-shot reasoning and structured event classification. The input video is temporally segmented into overlapping intervals $\mathcal{S}=\{s_{j}\}_{j=1}^{n}$ (e.g., 10 seconds each), where consecutive segments share partial temporal overlap to ensure continuous coverage of key activities and smooth transition detection. The 10 second interval duration was determined through pilot testing with 10 videos, balancing sufficient temporal context for meaningful event identification against computational efficiency and the model’s context window limitations. For each interval, the pipeline (Fig. 7) returns the following: the event type $e\in\mathcal{E}$, a natural language description $\mathcal{D}(e)$, an explanatory rationale $\mathcal{R}(e)$, a confidence level $l(e)\in\{\textit{high},\textit{medium},\textit{low}\}$, and representative frame(s) $\mathcal{K}(e)$ selected via keyframe extraction. A keyframe is selected as the frame with the highest count of distinct objects. This frame serves as the clearest snapshot of the interval. The model is instructed to state its confidence as one of three discrete levels, high, medium, or low, based on how certain it is about the inferred event type and description. This confidence level is passed unchanged to the front-end (e.g. shown as “High” in Fig. 3) and used as an ordinal indicator of reliability. This design follows recent work showing that large language models can express reasonably calibrated uncertainty in natural language without exposing logits or explicit probability scores (Lin et al., 2022; Kadavath et al., 2022).
@@ 5.2.2. Descriptor-based Search (DR5)
Objects are extracted from videos using BoxMOT (Broström, 2023), combining YOLO-based object detectors (Redmon et al., 2016) with DeepOCSort (Maggiolino et al., 2023) for multi-object tracking. LLAVA 1.6 (Liu et al., 2024a) verifies detection semantic correctness through binary questions, discarding negatives. For cars and persons, descriptor-based attributes are extracted through structured queries, with cropped first-appearance images stored for validation. Predicted attributes are re-evaluated on saved crops, and rejected outputs are placed in the “other” category.
@@ 7.1.1. Testbed environment
  • The EoIs Types (see Table 2) were informed by prior anomaly detection video datasets that define “anomaly” from a computer vision perspective as statistical rarity (Mahadevan et al., 2010), researcher-defined categories (Ramachandra and Jones, 2020), or labeled from online surveillance footage (Wu et al., 2020), valuable for benchmarking but not directly actionable in public safety practice. We drew EoIs from real campus crime logs (following Clery Act (Center, 2026)) on the foundation of existing anomaly datasets (Ma et al., 2015; Ramachandra and Jones, 2020; Qian et al., 2025; Wu et al., 2020; Wang et al., 2020a; Zhang et al., 2016; Acsintoae et al., 2022; Pranav et al., 2020; Lu et al., 2013; Rodrigues et al., 2020), then refined with frontline professionals to yield operationally meaningful, context-rich categories (e.g., “Incident Exposure”). This approach grounds “what to detect” in professionals’ operational realities, translating their priorities into actionable EoIs for future system and model design. However, our EoIs remain campus-centric, reflecting a limited jurisdictional perspective and excluding non-visual or hard-to-simulate incidents. We therefore position it not as a universal ontology, but as a reusable approach: MRVS-like systems should expose EoIs as configurable, locally governed “recipes” rather than fixed, one-size-fits-all classes. The EoI types serve as a reusable template that other jurisdictions can adapt, extend to additional modalities (e.g., audio, text), and use to drive MRVS-like systems and upstream workflows (e.g., dispatch (Liu et al., 2023; Terrell et al., 2004; Munro et al., 2025)).
@@ Acknowledgements.
This work was supported by the National Science Foundation (NSF2128867, NSF2350352), the Army Research Office (W911NF2320004, W911NF2520011), Google DeepMind (GDM), Clearpath Robotics, FrodoBots Lab, Raytheon Technologies (RTX), Tangenta, the Mason Innovation Exchange (MIX), and Walmart. We thank the George Mason University undergraduate students for their help with video dataset collection. We are grateful to the Northern Virginia police departments for their support and domain guidance, especially Detective Cooper Knight, Major Bull Hudson, Lieutenant Mohammed Tabibi, and Captain Andrew E. Hawkins, as well as other colleagues who provided feedback and guidance. We used generative AI tools for proofreading and editing under full human supervision, in line with the GAIDeT taxonomy (2025) (Suchikova et al., 2025). The authors are solely responsible for the final manuscript and outcomes.